# Ctrl-V: Higher Fidelity Video Generation with Bounding-Box Controlled Object Motion

### v2

@@ Abstract
With recent advances in video prediction, controllable video generation has been attracting more attention. Generating high fidelity videos according to simple and flexible conditioning is of particular interest. To this end, we propose a controllable video generation model using pixel level renderings of 2D or 3D bounding boxes as conditioning. In addition, we also create a bounding box predictor that, given the initial and ending frames’ bounding boxes, can predict up to 15 bounding boxes per frame for all the frames in a 25-frame clip. We perform experiments across 3 well-known AV video datasets: KITTI, Virtual-KITTI 2 and BDD100k.
Project page: https://oooolga.github.io/ctrl-v.github.io/.
Figure 1: Overview of Ctrl-V’s generation pipeline: Inputs (left): Our inputs include an initial frame, its corresponding bounding boxes and the final frame’s bounding boxes. Bounding-box prediction samples (middle): We illustrate three different sequences generated from our diffusion based bounding box motion prediction model. Videos sampled from our Box2Videodiffusion model  (right): Our Box2Video model conditions on the predicted bounding-box samples to produce final video clips.
@@ 1 Introduction
There has been rapid recent progress in controllable image generation with text-to-image diffusion models [1, 2]. Given conditioning such as sketches, depth maps, segmentation maps, skeleton poses, etc. one can now generate highly realistic images adhering to the given conditioning. However, to more accurately model the real world for applications like autonomous vehicle navigation and safety one needs to account for time.
It is widely recognized that most of the state-of-the-art, real-world autonomous vehicles systems are trained and/or tested in some manner in simulation. Current AV simulators generally rely on traditional hand-engineered computer graphics environments, as exemplified by the well-known (open source) CARLA simulator [3]. However, multiple experts in the field have envisioned a future where it is possible to construct simulators from generative models, potentially revolutionizing the way we train and test autonomous vehicles. Commercial grade neural sensor simulation approaches, like the one described by Yang et al., are starting to emerge publicly.
In this work, we focus upon the specific task of generating higher fidelity controllable video through developing a method that generates first person vehicle videos conditioned on predictions of 2D and 3D bounding boxes. Our method uses a diffusion approach to predict the motions and dynamics of objects in the form of sequences of bounding boxes, and generates video according to those bounding boxes. In our experiments, we verify that it is possible to generate video that adheres tightly to the desired bounding box motion conditioning, accurately depicting desired object movements. Additionally, it allows for supervisions on the appearance and disappearance of objects in a scene, including cars, pedestrians, bikers, and others. We call our approach Ctrl-V, a diffusion based bounding-box conditional video generation method that makes the following contributions and addresses multiple challenges associated with generating higher fidelity video from diffusion techniques:
@@
    2D-Bounding-Box and 3D-Bounding-Box Conditioning: Unlike our prior work, our approach extends beyond merely managing motion with 2D-bounding boxes; it also permits 3D-bounding-boxes to control image generation, providing an even more detailed level of control.
@@
    Bounding-box Motion Predictions with Diffusion: We present a novel diffusion based approach for predicting bounding-box trajectories based on their initial and final states.
@@
    Uninitialized Object Generation: Most bounding-box based generation only concerns with objects that are either present throughout the entire clip, or objects that are present at least in the first frame and persist for a while until the middle of the clip. New bounding boxes that appear after the first frame are often discarded or not accounted for [5]. In this work, we train our model to be sensitive to all bounding boxes whether they are present from the first frame and on, or only appear from the middle of the clip.
  4. 4.
    A New Benchmark for a New Problem Formulation: Given the novelty of our problem formulation, there is no existing standard way to evaluate models that seek to predict vehicle video with high fidelity. We therefore present a new benchmark consisting of a particular way of evaluating video generation models using the KITTI [6], Virtual KITTI 2 (vKITTI) [7] and the Berkeley Driving Dataset (BDD 100k) [8].
@@ 2 Related Work
Video latent diffusion models (VLDMs) extend latent image diffusion techniques [9] to video generation. Early VLDMs [10, 11, 12, 13, 14] shows temporally consistent frame generation and are tailored for text-prompted or image-prompted video generation. However, these models often struggle with complex scenes and lack the capability for precise local control.
@@
Conditional Video Diffusion techniques providing a certain degree of control. Methods like VideoCompose [15], Dreamix [16], Pix2Video [17], and DreamPose [18] propose various designs of novel adapters on top of VLDMs in order to incorporate different conditioning to achieve frame-level control. ControlNet Adapted Video Diffusion, on the other hand, achieve precise regional or pixel-level control in video generation by utilizing ControlNet [1] adapters within VLDM frameworks. Models such as Control-A-Video [19], Video ControlNet [20, 21], ControlVideo [22], and ReVideo [23] show that these adapters are highly adaptable to various types of conditioning, easy to train, and allow for more precise manipulation and enhanced accuracy in editing and creating video content.
@@
Motion Control with Bounding-Box Conditioning There are many strategies of control that have been explored in controllable video generation research. Notably, ControlVideo [22] utilizes a training-free strategy that employs pre-trained image LDMs and ControlNets to generate videos based on canny edge and depth maps. Control-A-Video [19] leverages a controllable video LDM that combines a pre-trained text-to-video model with ControlNet to manipulate videos using similar visual information. Video ControlNets [20, 21] uses optical flow information to enhance video generation, while ReVideo [23] depends on extracted video trajectories. DreamPose [18] injects pose sequence information into the initial noise. VideoComposer [15] uses an array of sketch, depth, mask, and motion vectors as conditioning.
@@
Many of these conditions, such as edge, depth, and optical flow maps, are costly to produce and lack the flexibility needed for customization. Bounding box emerge as a conditioning that are easily customizable and can be edited into different shape, size, locations and classes efficiently. To the best of our knowledge, five other research projects are currently exploring the use of bounding boxes for motion control in video generation. However, it is important to note that our work is distinct from these in several critical respects.
@@
TrailBlazer [24] and Peekaboo [25] are different training-free approaches that employ attention map adjustments to direct the model in generating a particular object within a defined region. FACTOR [26] augmented the transformer-based generation model, Phenaki [27], by integrating a box control module. TrailBlazer, Peekaboo and FACTOR necessitate textual descriptions for individual boxes, thus lacking direct visual grounding.
@@
Our task setup aligns closely with that of Boximator[5] and TrackDiffusion [28], in that we also do not require text descriptions for individual boxes, relying solely on bounding box inputs. However, our approaches diverge significantly. Boximator and TrackDiffusion enhance their models by introducing new self-attention layers to 3D U-Net blocks. These layers incorporate additional conditional information, such as box coordinates and object IDs, into the pretrained VLDM model. Their bounding-box information is processed using a Fourier embedder [29], which is then passed through multi-layer perceptron layers to encode. In contrast, our approach uses ControlNet and does not involve training additional encoding layers or utilizing Fourier embedder to handle the bounding-box information. Moreover, Boximator introduces a self-tracking technique to ensure adherence to bounding-box in generated outputs, a technique also adopted by TrackDiffusion. This enables the network to learn the object tracking task alongside video generation, but requires a two-stage training process: one with target bounding boxes in frames, and another with the boxes removed. They demonstrate that without this technique, the model’s performance markedly declines. Conversely, our model achieves alignment with the bounding-box conditions without additional training.
@@
Vehicle Oriented Generative Models DriveDreamer [30] presents noteworthy contribution from autonomous driving domain. It takes an action-based approach to video simulation. It also makes use of bounding boxes and generate actions along with a video rendering. Within the DriveDreamer framework, Fourier embeddings [29] are also employed to encode bounding-box information, and CLIP embeddings [31] are used for box categorization. They focus on generating multiple camera views and do not condition on bounding box sequences, so cannot be directly compared with our problem setting. In contrast, the DriveGAN work of [32] aims to learn a GAN based driving environment in pixel-space, complete with actions and an implicit model of dynamics encoded using the latent space of a VAE. While driving oriented, the approach does not focus on controlling the generation of vehicle video that respects well-defined object trajectories with high fidelity.
@@ 3 Our Method: Ctrl-V
In this section, we outline our simple yet robust method for creating video clips that strictly adhere to specified frame-level bounding-box guidelines. Our approach has two steps, involving the training of two distinct models: 1) the fine-tuning of a Video Latent Diffusion Model (VLDM) to predict bounding-box frames (Section 3.3) – which are images of only bounding boxes; 2) the training of our Box2Video to enable precise frame-level conditioning (Section 3.4).
@@ 3.1 Preliminaries
Stable Video Diffusion (SVD) [10, 11] is a cutting-edge video generation model that excels in converting still images into dynamic video sequences. The pretrained SVD model employed in this project has been pretrained using the Euler discrete noise scheduling method outlined by Karras et al. [33]. The objective of the SVD model is to generate a video sequence ${\bm{f}}=[{\bm{f}}^{(0)},\ldots,{\bm{f}}^{(N)}]$ starting from an initial frame ${\bm{f}}^{(0)}$, where $N$ represents the sequence length.
The SVD model utilizes an image autoencoder to encode and decode each frame, converting them between pixel space and latent space. Mathematically, this process is represented as $\mathcal{D}\big{(}\mathcal{E}({\bm{f}})\big{)}=\mathcal{D}({\bm{z}})\approx{\bm{f}}$. In SVD, the diffusion process occurs within the latent space. This process entails gradually introducing noise to the latent representations. Subsequently, the denoiser network is responsible for removing this added noise from the noisy latent representations. The denoiser network, trained under the Euler discrete noise scheduling schema, can be described by the formulation:
${\mathbb{D}}_{\theta}({\bm{z}};\sigma_{t})=\lambda_{\text{skip}}(\sigma_{t}){\bm{z}}+\lambda_{\text{out}}(\sigma_{t}){\mathbb{U}}_{\theta}\big{(}\lambda_{\text{in}}(\sigma_{t}){\bm{z}},{\bm{z}}^{(0)}_{\text{pad}},{\bm{c}}^{(0)};\lambda_{\text{noise}}(\sigma_{t})\big{)}$ (1)
Here $\lambda_{\text{skip}}$, $\lambda_{\text{out}}$ and $\lambda_{\text{in}}$ are scaling functions, the mathematical definitions of which can be found in the work of Karras et al. [33]. ${\mathbb{U}}_{\theta}$ represents the 3D U-Net [34], whose parameters $\theta$ are optimized for the denoising task during training.
To gain a clearer understanding of the U-Net inputs, we eliminate the scaling terms and reparameterize its input as follows:
${\mathbb{U}}_{\theta}\big{(}\hat{{\bm{z}}}_{t},{\bm{z}}_{\text{pad}}^{(0)},{\bm{c}}^{(0)},t\big{)}$ (2)
where $\hat{{\bm{z}}}_{t}\in\mathbb{R}^{N,C^{\prime},H^{\prime},W^{\prime}}$ denotes the latent representation of frames corrupted by noise at the current noise level $t$; ${\bm{z}}^{(0)}\in\mathbb{R}^{1,C^{\prime},H^{\prime},W^{\prime}}$ represents the latent representation of the initial frame, and ${\bm{c}}^{(0)}$ represents the encoding of the initial frame image token using CLIP [31]. These parameters are introduced into the U-Net through distinct layers. Initially, at the input layer, ${\bm{z}}^{(0)}$ undergoes repeated padding along the frame’s time dimension to match the dimensions of the $\hat{{\bm{z}}}_{t}$ tensor, namely ${\bm{z}}_{\text{pad}}^{(0)}$. Subsequently, the ${\bm{z}}_{\text{pad}}^{(0)}$ is concatenated with $\hat{{\bm{z}}}_{t}$ along the channel dimension and passed into the network’s input layer. Meanwhile, the noise level indicator $t$ is transformed into a time-step embedding vector and is subsequently fed into each U-Net block along with ${\bm{c}}^{(0)}$ via cross-attention conditioning mechanism [9].
3.2 Preprocessing: Bounding-Box Frame Preparation
Instead of resorting to Fourier transform or other sequential models to encode bounding-box information in string format, we have chosen to represent the bounding boxes through image rendering. We created a method to make bounding-box plots which have significant benefits. These plots not only show the precise locations of objects within a scene but also effectively display different kinds of information, like track IDs and object types, using different shapes and colors. We also implement a method to render trajectory frames, similar to the bounding-box frames, but without the pixel-space size information of the objects. These trajectory frames only contain mid-point location information. The procedure for creating bounding-box/trajectory plots is detailed in Appendix A.1. Additionally, we utilize the image-encoder ($\mathcal{E}$) to encoder the bounding-box frames, and its performance is discussed in Appendix A.2.
Figure 2: The overview of Ctrl-V: (left) BBox Predictor and (right) Box2Video.
3.3 Ctrl-V: BBox Predictor
Our bounding-box predictor aims to forecast object positions across intermediate frames, utilizing the initial frame, initial bounding-box frame(s) and the final bounding-box/trajectory frame as references. Our approach involves adjusting the SVD network and leveraging it to generate bounding-box frames. Specifically, we modify the inputs of the U-Net, namely ${\bm{z}}_{t}$, ${\bm{z}}_{\text{pad}}^{(0)}$, and ${\bm{c}}^{(0)}$ (refer to Equation 2), to achieve this objective.
Initially, we pass the bounding-box frames through the image encoder ($\mathcal{E}$) to obtain their latent representations, denoted as $\mathcal{E}({\bm{f}}_{\text{bbox}})={\bm{b}}\in\mathbb{R}^{N,C^{\prime},H^{\prime},W^{\prime}}$. Furthermore, we feed the initial frame into both the image encoder and the CLIP encoder, resulting in ${\bm{z}}^{(0)}\in\mathbb{R}^{1\times C^{\prime}\times H^{\prime}\times W^{\prime}}$ and ${\bm{c}}^{(0)}$ respectively.
In alignment with the original SVD configuration, we replicate the padding procedure from ${\bm{z}}^{(0)}\in\mathbb{R}^{1\times C^{\prime}\times H^{\prime}\times W^{\prime}}$ to ${\bm{z}}^{(0)}_{\text{pad}}\in\mathbb{R}^{N\times C^{\prime}\times H^{\prime}\times W^{\prime}}$. However, in this scenario, we replace the first and last entries of ${\bm{z}}^{(0)}_{\text{pad}}$ with ${\bm{b}}^{(0)}$ and ${\bm{b}}^{(N-1)}$ respectively. Moreover, depending on the desired number of initial bounding-box frame(s) for conditioning, we also experiment with substituting various numbers of the ${\bm{z}}^{(0)}$ frames with bounding-box frames. In Equation 3.3, $\tilde{{\bm{b}}}$ denotes the modified tensor.
Furthermore, we replaced the $\hat{{\bm{z}}}_{t}$ in Equation 2 with $\hat{{\bm{b}}}_{t}$, where noise at level $t$ is incorporated into ${\bm{b}}$. The primary objective of the denoiser network is to estimate the amount of noise introduced to ${\bm{b}}$.
Overall, the input to the bounding-box predictor’s U-Net is represented as ${\mathbb{U}}^{\text{bbox}}_{\omega}\big{(}\hat{{\bm{b}}}_{t},\tilde{{\bm{b}}},{\bm{c}}^{(0)},t\big{)}$.

### v3

# Ctrl-V: High Fidelity Video Generation with Bounding-Box Controlled Object Motion

@@ Abstract
Controllable video generation has attracted significant attention, largely due to advances in video diffusion models. In domains such as autonomous driving, it is essential to develop highly accurate predictions for object motions. This paper tackles a crucial challenge of how to exert precise control over object motion for realistic video synthesis. To accomplish this, we 1) control object movements using bounding boxes and extend this control to the renderings of 2D or 3D boxes in pixel space, 2) employ a distinct, specialized model to forecast the trajectories of object bounding boxes based on their previous and, if desired, future positions, and 3) adapt and enhance a separate video diffusion network to create video content based on these high quality trajectory forecasts. Our method, Ctrl-V, leverages modified and fine-tuned Stable Video Diffusion (SVD) models to solve both trajectory and video generation. Extensive experiments conducted on the KITTI, Virtual-KITTI 2, BDD100k, and nuScenes datasets validate the effectiveness of our approach in producing realistic and controllable video generation.
Figure 1: Overview of Ctrl-V’s generation pipeline: (Left) inputs: Our inputs include an initial frame, its corresponding bounding box image and the final frame’s bounding box image. (Middle) generated bounding box trajectories: We demonstrate three distinct possible trajectory sequences produced by our diffusion-based bounding box motion generation model – BBox Generator. (Right) generated video clips: Our Box2Video model conditions on the generated bounding box trajectory frames to produce the final video clips.
@@ 1 Introduction
Recent advances in controllable image generation have enabled the creation of highly realistic images from various conditioning inputs, including points, bounding boxes, scribbles, segmentation maps, and skeleton poses. Yet, translating this control to video generation is markedly more challenging due to the added temporal dimension. Incorporating time dynamics into diffusion models significantly complicates controllable video generation, as it requires accounting for object interactions, physical consistency, and coherent motion across frames.
Numerous recent studies have examined different forms of controllability for video generation. Researchers have used an array of methods for control, including conditioning on information such as canny edge and depth maps (Zhang et al. (2023b)), similar visual information (Chen et al. (2023)), optical flow (Hu & Xu (2023)), and pose sequences (Karras et al. (2023)). These control inputs are often expensive to produce, especially when sequences of them are required in order to condition a video. Models that use accessible conditioning such as bounding boxes require additional input such as text to help with the generation process (Wang et al. (2024)). A controllable video generation model with an accessible and simple mode of control is greatly desired.
In this work, we focus on creating such a model. Specifically, we aim to generate higher fidelity videos controlled by, at the minimum, the beginning and ending positions of 2D and 3D bounding boxes without the help of other modes of control. Our two-part method includes a diffusion-based model that generates the motions and dynamics of objects in the form of bounding box videos (2D images of the bounding boxes evolving over time), and a generative model of videos according to those bounding box videos. To this end, we choose to train and test our model on driving datasets as they contain challenging scenes rich with different types of bounding boxes as well as complex movement and irregular appearing and disappearing objects. In our experiments, we show that our model generates videos that adhere tightly to the desired bounding box motion conditioning, accurately depicting desired object movements. Additionally, through our novel pixel-level bounding box generator and conditioning, our method robustly handles the appearance and disappearance of different objects in a scene, including cars, pedestrians, bikers, and others.
In this paper, we present Ctrl-V, a diffusion-based bounding box conditional video generation method that addresses multiple challenges and makes the following contributions to generate higher-fidelity videos using diffusion techniques. Our contributions can be enumerated as follows:
@@
    A Novel Diffusion Approach for Predicting Bounding Box Trajectories: Our approach generates video frames with 2D or 3D bounding box trajectories at the pixel-level based on their initial and final states, and the first frame of RGB video. Our results show that our proposed approach yields higher quality bounding box trajectory predictions than a recently proposed state of the art method (see in Table 1).
@@
    2D and 3D Bounding Box Conditioned Video Generation with Diffusion: We present a method that allows video generation by conditioning on 2D or 3D bounding box trajectories which allows fine-grained control over the generated videos. Our results (in Table 2) indicate that this approach can generate video that is dramatically better than a state-of-the-art Stable Video Diffusion baseline fine-tuned without this conditioning mechanism across four commonly used quality metrics. Our approach further improves upon prior work by enabling the following capabilities: a. Multi-subject Generation: Synthesizing multiple subjects in videos poses significant challenges, requiring coherent object placement across frames, particularly during interactions. Existing models typically demonstrate capabilities with up to three subjects in online demo clips. Recent advances include: Boximator (Wang et al., 2024) (which can synthesize up to 8 subjects) and FACTOR (Li et al., 2024) (up to 12 subjects). Our method enables synthesis of scenes with any number of objects, limited only by the number of clearly renderable bounding boxes per frame; b. Uninitialized Object Generation: Most bounding box-based generation methods focus solely on objects that either remain present throughout the entire clip or appear in the first frame and persist until at least the middle of the clip. They typically overlook bounding boxes that appear only after the first frame (Wang et al., 2024). In this work, we train our model to be sensitive to all bounding boxes, whether they are present from the first frame or appear from the middle of the clip.
@@
    A New Benchmark for a New Problem Formulation: Given the novelty of our problem formulation, there is no existing standard way to evaluate models that seek to predict vehicle video with high fidelity. We therefore present a new benchmark consisting of a particular way of evaluating video generation models using the KITTI (Geiger et al., 2013), Virtual KITTI 2 (vKITTI) (Cabon et al., 2020), the Berkeley Driving Dataset (BDD 100k) (Yu et al., 2020) and nuScenes (Caesar et al., 2019).
@@ 2 Related Work
Video latent diffusion models (VLDMs) extend latent image diffusion techniques (Rombach et al., 2022) to video generation. Early VLDMs (Blattmann et al., 2023b; a; He et al., 2023; Zeng et al., 2023; Wu et al., 2023) shows temporally consistent frame generation and are tailored for text-prompted or image-prompted video generation. However, these models often struggle with complex scenes and lack the capability for precise local control.
@@
Conditional Video Diffusion techniques providing a certain degree of control. Methods like VideoCompose (Wang et al., 2023a), Dreamix (Molad et al., 2023), Pix2Video (Ceylan et al., 2023), and DreamPose (Karras et al., 2023) propose various designs of novel adapters on top of VLDMs in order to incorporate different conditioning to achieve frame-level control. ControlNet Adapted Video Diffusion, on the other hand, achieve precise regional or pixel-level control in video generation by utilizing ControlNet (Zhang et al., 2023a) adapters within VLDM frameworks. Models such as Control-A-Video (Chen et al., 2023), Video ControlNet (Hu & Xu, 2023; Chu et al., 2023), ControlVideo (Zhang et al., 2023b), and ReVideo (Mou et al., 2024) show that these adapters are highly adaptable to various types of conditioning, easy to train, and allow for more precise manipulation and enhanced accuracy in editing and creating video content.
@@
Motion Control with bounding box Conditioning There are many strategies of control that have been explored in controllable video generation research. Notably, ControlVideo (Zhang et al., 2023b) utilizes a training-free strategy that employs pre-trained image LDMs and ControlNets to generate videos based on canny edge and depth maps. Control-A-Video (Chen et al., 2023) leverages a controllable video LDM that combines a pre-trained text-to-video model with ControlNet to manipulate videos using similar visual information. Video ControlNets (Hu & Xu, 2023; Chu et al., 2023) uses optical flow information to enhance video generation, while ReVideo (Mou et al., 2024) depends on extracted video trajectories. DreamPose (Karras et al., 2023) injects pose sequence information into the initial noise. VideoComposer (Wang et al., 2023a) uses an array of sketch, depth, mask, and motion vectors as conditioning.
@@
Many of these conditions, such as edge, depth, and optical flow maps, are costly to produce and lack the flexibility needed for customization. Bounding boxes emerge as a conditioning that are easily customizable and can be edited into different shape, size, locations and classes efficiently. To the best of our knowledge, six other research projects are currently exploring the use of bounding boxes for motion control in video generation. However, it is important to note that our work is distinct from these in several critical respects.
@@
Direct-A-Video, TrailBlazer (Ma et al., 2024) and Peekaboo (Jain et al., 2024) are different training-free approaches that employ attention map adjustments to direct the model in generating a particular object within a defined region. Direct-A-Video, in particular, is a text-to-video model that learns to control camera motion during training and then adopts a training-free approach to manipulate object movements using bounding boxes. FACTOR (Huang et al., 2023) augmented the transformer-based generation model, Phenaki (Villegas et al., 2022), by integrating a box control module. TrailBlazer, Peekaboo and FACTOR necessitate textual descriptions for individual boxes, thus lacking direct visual grounding.
@@
Our task setup shares mild similarities with Boximator(Wang et al., 2024) and TrackDiffusion(Fischer et al., 2023) because we also utilize bounding box conditioning during training without relying on text descriptions for individual boxes. However, our approach diverges from these text-to-video models, as our primary focus is on generating realistic videos conditioned only on a couple frames of bounding boxes, whereas Boximator and TrackDiffusion are designed to be conditioned on text information as they both are text-to-video models. Boximator and TrackDiffusion enhance their models by introducing new self-attention layers to 3D U-Net blocks. These layers incorporate additional conditional information, such as box coordinates and object IDs, into the pretrained VLDM model. Their bounding box information is processed using a Fourier embedder (Mildenhall et al., 2020), which is then passed through multi-layer perceptron layers to encode. In contrast, our approach uses ControlNet and does not involve training additional encoding layers or utilizing Fourier embedder to handle the bounding box information. Moreover, Boximator introduces a self-tracking technique to ensure adherence to the bounding boxes in generated outputs, a technique also adopted by TrackDiffusion. This enables the network to learn the object tracking task alongside video generation, but requires a two-stage training process: one with target bounding boxes in frames, and another with the boxes removed. They demonstrate that without this technique, the model’s performance markedly declines. Conversely, our model achieves alignment with the bounding box conditions without additional training.
@@
Vehicle Oriented Generative Models DriveDreamer (Wang et al., 2023b) presents noteworthy contribution from autonomous driving domain. It takes an action-based approach to video simulation. It also makes use of bounding boxes and generate actions along with a video rendering. Within the DriveDreamer framework, Fourier embeddings (Mildenhall et al., 2020) are also employed to encode bounding box information, and CLIP embeddings (Radford et al., 2021) are used for box categorization. They focus on generating multiple camera views and do not condition on bounding box sequences, so cannot be directly compared with our problem setting. In contrast, the DriveGAN work of Kim et al. (2021) aims to learn a GAN based driving environment in pixel-space, complete with actions and an implicit model of dynamics encoded using the latent space of a VAE. While driving oriented, the approach does not focus on controlling the generation of vehicle video that respects well-defined object trajectories with high fidelity.
@@ 3.1 Preliminaries
We begin here with an overview of the Stable Video Diffusion (SVD) model (Blattmann et al., 2023a), due to its importance in our approach. SVD is a diffusion based image-to-video (I2V) model performed in latent embedding space Blattmann et al. (2023b). Using an image ${\bm{f}}^{(0)}$ as initial condition, SVD is able to extend that single frame into a video ${\bm{f}}=[{\bm{f}}^{(0)},\ldots,{\bm{f}}^{(N)}]$ where $N$ is the length of the sequence. Notably, SVD operates in latent space, where the diffusion and denoising process act upon the latents ${\bm{z}}$ of the video ${\bm{f}}$. Here, SVD employs an image encoder ($\mathcal{E}$) and an image decoder ($\mathcal{D}$) to translate each frame into and out of latent space: $\mathcal{D}\big{(}\mathcal{E}({\bm{f}}^{(i)})\big{)}=\mathcal{D}({\bm{z}}^{(i)})\approx{\bm{f}}^{(i)}$. At each diffusion step, SVD progressively introduces noise into the latent representations. In this work, the amount of noise is dictated by Euler discrete noise scheduling method (EDM) introduced in Karras et al. (2022). A UNet based denoiser network within the SVD is used to predict this noise in order to recover the original latent representations. The UNet, ${\mathbb{U}}_{\theta}$, is parameterized as:
${\mathbb{U}}_{\theta}\big{(}\hat{{\bm{z}}}_{t},{\bm{z}}_{\text{pad}}^{(0)},{\bm{c}}^{(0)},t\big{)},$ (1)
  • $\hat{{\bm{z}}}_{t}\in\mathbb{R}^{N\times C^{\prime}\times H^{\prime}\times W^{\prime}}$: latent representation of frames corrupted by noise at noise level $t$.
  • ${\bm{z}}^{(0)}\in\mathbb{R}^{1\times C^{\prime}\times H^{\prime}\times W^{\prime}}$: latent representation of the initial frame.
  • ${\bm{z}}_{\text{pad}}^{(0)}\in\mathbb{R}^{N\times C^{\prime}\times H^{\prime}\times W^{\prime}}$: Padded ${\bm{z}}^{(0)}$ by repeating itself along the first dimension N times.
  • ${\bm{c}}^{(0)}$: CLIP encoding (Radford et al., 2021) of the initial frame.
The full denoiser network, ${\mathbb{D}}_{\theta}$, with an EDM noise scheduler, is formulated as
${\mathbb{D}}_{\theta}({\bm{z}};{\bm{c}}^{(0)},\sigma_{t})=\lambda_{\text{skip}}(\sigma_{t}){\bm{z}}+\lambda_{\text{out}}(\sigma_{t}){\mathbb{U}}_{\theta}\big{(}\lambda_{\text{in}}(\sigma_{t}){\bm{z}},{\bm{z}}^{(0)}_{\text{pad}},{\bm{c}}^{(0)};\lambda_{\text{noise}}(\sigma_{t})\big{)}$ (2)
Here $\lambda_{\text{skip}}$, $\lambda_{\text{out}}$, $\lambda_{\text{in}}$ and $\lambda_{\text{noise}}$ denote scaling functions, while $\sigma_{t}$ represents the computed noise at level $t$. The precise mathematical definitions of these terms are detailed in Appendix B. Note that 3D UNet ${\mathbb{U}}_{\theta}$ in Equation 1, is a re-parameterized version of the one in Equation 2 (Ronneberger et al., 2015). The scaling terms are absorbed and the inputs are simplified for clarity. In the following sections, we follow the re-parameterized version in Equation 1 when refering to the UNets in our model.
3.2 Overview of our Method: Ctrl-V
Our controllable video generation method is illustrated in Figure 2. It consists of two sequential steps:
  1. 1.
    First, we generate bounding box frames using our diffusion based bounding box predictor network, the BBox Generator, which is shown on the left side of Figure 2. These frames contain only bounding boxes. They make up a video of moving (or stationary) bounding boxes and it serve as the “skeleton” for the generated video.
  2. 2.
    Then, we generate a video using our video generator network, Box2Video, shown on the right side of Figure 2, where the bounding boxes frames act as the control signal. The bounding boxes in each frame determine the objects generated in the corresponding frames of the video.
BBox Generator and Box2Video each utilizes a modified SVD backbone – illustrated by the SVD backbone in Figure 2. These backbones are adapted to their respective generation tasks. Details of each model are presented in their individual sections: BBox Generator – Section  3.3 and Box2Video – Section  3.4.
[…]
3.3 Ctrl-V: BBox Generator
The BBox Generator shown on the left in Figure 2 aims to predict object bounding boxes across all video frames using an SVD backbone. The four inputs to the model are $\hat{{\bm{b}}}_{t}$, ${\bm{b}}^{(0)}$, ${\bm{b}}^{(N-1)}$, ${\bm{z}}^{(0)}$, where: $\hat{{\bm{b}}}_{t}$ is the encoded “video” of bounding boxes with $t$ levels of noise added; ${\bm{b}}^{(0)}$ is the encoded initial bounding box frame(s); ${\bm{b}}^{(N-1)}$ is the encoded final bounding box frame; ${\bm{z}}^{(0)}$ is the encoded initial video frame. During training, the model learns to predict the noise added in $\hat{{\bm{b}}}_{t}$ according to the EDM noise scheduler. The model recovers the original ${\bm{b}}$ from its noisy version $\hat{{\bm{b}}}_{t}$ by calculating the noise with UNet outputs and eliminating the noise through scaling functions. We opt to abstract this detail in the model diagram for readability.
In practice, the four inputs are transformed and concatenated into a vector format accepted by the UNet adapter within the SVD backbone. Specifically, as shown in Figure 2, ${\bm{z}}^{(0)}\in\mathbb{R}^{1\times C^{\prime}\times H^{\prime}\times W^{\prime}}$ is replicated along the first dimension, and its front and end (in the first dimension) are replaced by ${\bm{b}}^{(0)}$, ${\bm{b}}^{(N-1)}$ respectively. This forms ${\bm{z}}_{\text{pad}}^{(0)}=\text{concat}({\bm{b}}^{(0)},{\bm{z}}^{(0)},...,{\bm{z}}^{(0)},{\bm{b}}^{(N-1)})\in\mathbb{R}^{N\times C^{\prime}\times H^{\prime}\times W^{\prime}}$. The noise-added encoding of bounding box video $\hat{{\bm{b}}}_{t}$ is then concatenated with ${\bm{z}}_{\text{pad}}^{(0)}$ to form the final input to the UNet adapter. The network incorporates additional conditioning inputs, including a CLIP-encoded embedding of the initial frame ${\bm{c}}^{(0)}$ and a noise-level embedding $t$. These embeddings are individually integrated into every sub-block of the U-Net through a self-attention mechanism.
Representing Bounding Boxes in Pixel Space
An important element of Ctrl-V is our design choice of rendering bounding boxes in pixel space. The manner in which bounding box information is provided as a control signal to the video generator is important. For example, prior work such as Boximator (Wang et al., 2024) represents bounding boxes as a Fourier transformed concatenated vector of their raw coordinates, ID and other information. In contrast, in our work we choose to render bounding boxes into frames while maintaining minimal loss of meta information. Importantly, we also encode information such as track ID, object type, and orientation for each bounding box using a combination of visual attributes, including border color, fill color, and markings. Specifically, the track ID represents a unique identifier for each tracked object across frames, the object type specifies the category of the object (e.g., car, pedestrian), and the orientation indicates the direction the object is facing. Further details about how these bounding box frames are rendered can be found in Appendix C.1. Crucially, our approach allows us to leverage the highly effective ControlNet approach to provide pixel-level guidance to influence diffusion generated imagery.